# LLMs’ Suitability for Network Security: A Case Study of STRIDE Threat Modeling Author’s draft for soliciting feedback - May 6, 2025

### v1

@@ Abstract
Artificial Intelligence (AI) is expected to be an integral part of next-generation AI-native 6G networks. With the prevalence of AI, researchers have identified numerous use cases of AI in network security. However, there are almost nonexistent studies that analyze the suitability of Large Language Models (LLMs) in network security. To fill this gap, we examine the suitability of LLMs in network security, particularly with the case study of STRIDE threat modeling. We utilize four prompting techniques with five LLMs to perform STRIDE classification of 5G threats. From our evaluation results, we point out key findings and detailed insights along with the explanation of the possible underlying factors influencing the behavior of LLMs in the modeling of certain threats. The numerical results and the insights support the necessity for adjusting and fine-tuning LLMs for network security use cases.
Index Terms:
Large Language Model (LLM), STRIDE, threat modeling, suitability of LLMI Introduction
@@
We provide detailed insights based on the evaluation results of LLM-based STRIDE classification. We provide detailed discussions on potential underlying factors that influence the behavior of LLMs in modeling certain threats, including: incorrect threat perspective, failure to identify second-order threats, and insights on Few-Shot (FS) prompting positively impacting performance.
@@ II Motivation and Examination of Related Work
LLM for Networking: The white paper by Shahid et al. [4] presents the concept of “Large Telecom Model (LTM)” for use cases of telecom networks. Some of the potential use cases they mention include the use of LTM at the network edge, LTM for network optimization, using LTM for network automation tasks, and performing intent-based management using LTMs, etc. Similarly, Zhou et al. [5] identify four different areas of telecom networks that may benefit from LLMs. These areas include generation, optimization, classification, and prediction problems in telecom networks. Wu et al. [11] present the NetLLM method to utilize LLMs for three different networking problems, namely, “viewport prediction”, “adaptive bitrate streaming”, and “cluster job scheduling”. The authors extensively evaluate the performance of their proposed framework within these problems.
LLM for Security: Ferrag et al. [6] propose multiple applications of LLMs in cybersecurity, including detection and analysis of threats, incident response, automation of security tasks, verification of security protocols, etc. The work by Guthula et al. [12] proposes a foundation model for security that takes into account the distinct nature of network traffic. The aim of the authors is to consider the general applicability of the model.
Sędkowski [13] studies the efficacy of LLMs in recognizing potential threats in the network and recommending countermeasures. Their methodology includes using Nmap reports to classify threats using STRIDE threat modeling with three different LLMs. The author concludes that the performance of LLMs in threat detection is comparable to humans. The scope of their work is focused on testing the application of AI in threat modeling, as opposed to our aim of studying the suitability of LLMs for network security.
Yang et al. [14] present “ThreatModeling-LLM”, which is an LLM-based method to perform threat modeling of the banking system using the STRIDE model. Their approach includes various steps to improve the performance of the LLMs. Their scope is limited to banking systems. The main objective of their work is to improve and automate threat modeling using LLMs, instead of investigating the suitability of LLMs for network security.
Motivation: After examining the related works, we identify that the suitability of the potential use cases of LLMs needs to be investigated. However, to the best of our knowledge, there are almost nonexistent or very scarce existing studies that are actually testing the suitability of LLMs for these use cases. This motivated us to study the suitability of the LLMs in telecom network security using a case study of STRIDE threat modeling.
@@ III Evaluation Methodology for A Case Study of STRIDE Threat Modeling
Our evaluation methodology is shown in Figure 1. We initially select multiple threats and vulnerabilities in the 5G network along with their baseline STRIDE classifications from the published literature and standards. Then, we use various prompting techniques to perform the LLM-based STRIDE classification of the selected threats. Finally, we compare the STRIDE classification by LLMs to the baseline and evaluate the results. We will explain each step in the following.
@@ III-A Selected 5G Threats and Vulnerabilities
To carry out the evaluation, we select six threats and vulnerabilities on 5G interfaces from our previous research work [10] along with their STRIDE classifications as a baseline; this step is shown with a blue block in Figure 1. As we already mentioned in our previous work [10], we want to clarify that the baseline STRIDE classification of the threats may not be unique. We randomly select these attacks from the published literature and standards and ensure that the selected threats cover the six STRIDE categories. Three of the selected threats are on the N1 interface, because N1 is exposed to the Radio Access Network (RAN) and it faces the largest number of threats [10]. In our earlier work [10], we comprehensively explored and identified the threats and vulnerabilities on the critical 5G network interfaces and categorized them based on the STRIDE threat model. The selected threats, along with their description, are explained below. We use the same threat names in the first row of Table I:
@@ III-B Large-Language Models (LLMs)
We select the following Large Language Models (LLMs)to perform this evaluation: Sonar by Perplexity [19], GPT-4o by OpenAI [20], Claude 3.7 Sonnet by Anthropic [21], Grok-2 by xAI [22], and Gemini 2.5 Pro by Google [23]. We use these LLMs through the pro version of the Perplexity AI platform that we have access to through our University [24]. We perform the STRIDE classification of the six selected threats using these LLMs in order to evaluate the suitability of the LLMs for network security. Figure 1 shows the LLM-based STRIDE classification methodology with green blocks.
@@ III-C LLM Prompts
We replace the ‘[NAME_OF_THREAT]’ in these prompts with the name of a specific threat and provide the prompt to the LLM. Then, we record the STRIDE classification of a threat provided by the LLM. We repeat this step for all the six selected threats (see steps in green blocks in Figure 1).
@@ IV-A LLM-based STRIDE Classification
In this section, we provide the results of evaluating the case study of STRIDE threat modeling of 5G threats. The results of LLM-based STRIDE classification of the six 5G threats are shown in Table I. The first column of the table includes the prompting techniques we employ in our evaluation, while the second column shows the LLM models we select. The rest of the columns in this table present the baseline STRIDE classification along with the results of LLM-based STRIDE classification of the 5G threats. In this table, the white cells with a dot represent a positive baseline value, while an empty white cell represents a negative baseline value. According to this baseline, we categorize the LLM-based STRIDE classifications as True Positive (TP) (dark green cell with a dot ), True Negative (TN) (empty green cell), False Positive (FP) (yellow cell with a dot ), and False Negative (FN) (empty red cell). The coloring scheme represents that the greens (TP and TN) are correct classifications. Yellow (FP) is an incorrect classification, but it is not the worst outcome (over-predicting positive), and red (FN) is an incorrect classification and represents the worst outcome (under-predicting positive). In the following, we present the main insights and observations from these results.
@@ IV-D Performance of LLMs in STRIDE Classification
The results show that LLMs’ performance in our experiments is mostly comparable. Figure 4 illustrates the performance of the LLMs in terms of accuracy, precision, recall, and F1 score using a heatmap chart. This result is averaged across all threats and all prompting techniques for a specific LLM. We observe that GPT-4o, Claude 3.7 Sonnet, and Grok-2 show ‘relatively’ higher accuracy and precision but lower recall compared to the other two LLMs. On the other hand, Sonar and Gemini 2.5 Pro achieved lower precision and higher recall in comparison. The F1 score indicates that the performance of all the LLMs is comparable for this case study. The maximum accuracy achieved is , which highlights that there are opportunities for improvement across all LLMs, perhaps by fine-tuning for the specific application of STRIDE threat modeling in 5G networks.
We note that several articles highlight the prominent challenges and issues with the LLMs [27, 28]. The most relevant challenges to threat modeling include incorrect predictions and LLM hallucinations, which may result in disregarding required countermeasures or implementing unnecessary security measures. Secondly, the adaptability of LLMs for telecom-specific threats and vulnerabilities. Thirdly, improving LLM inference speed in networks to enable rapid threat modeling of the detected threats.
@@ V Discussion and Conclusion
In this work, we explore and investigate the suitability of LLMs for network threat modeling. To perform this analysis, we select the case study of STRIDE threat modeling to perform LLM-based classification of 5G threats according to the STRIDE threat model. We observe from our evaluation that providing examples to the LLMs using FS prompting improves their performance. We further notice that LLMs may not always consider the threat classifications as a result of the second-order effect. This will limit the threat identification and may eventually result in not identifying all the possible risks associated with a threat. We hope these insights and results are the starting points to encourage research into fine-tuning LLMs on telecom-specific datasets and to enhance their performance in network security tasks. This is particularly important for future ‘AI-native’ networks, where AI needs to detect and identify threats autonomously and with the highest accuracy.

### v2

@@ Abstract
Artificial Intelligence (AI) is expected to be an integral part of next-generation AI-native 6G networks. With the prevalence of AI, researchers have identified numerous use cases of AI in network security. However, there are very few studies that analyze the suitability of Large Language Models (LLMs) in network security. To fill this gap, we examine the suitability of LLMs in network security, particularly with the case study of STRIDE threat modeling. We utilize four prompting techniques with five LLMs to perform STRIDE classification of 5G threats. From our evaluation results, we point out key findings and detailed insights along with the explanation of the possible underlying factors influencing the behavior of LLMs in the modeling of certain threats. The numerical results and the insights support the necessity for adjusting and fine-tuning LLMs for network security use cases.
I Introduction
@@
We provide detailed insights based on the evaluation results of LLM-based STRIDE classification. We present detailed discussions on potential underlying factors that influence the behavior of LLMs in modeling certain threats, including incorrect threat perspective, failure to identify second-order threats, and insights on Few-Shot (FS) prompting positively impacting performance.
@@
It should be noted for clarity that our study focuses on classifications and predictions using LLMs in the context of STRIDE threat modeling. Therefore, other performance metrics such as inference speed, adaptability, or scalability of LLMs are outside the scope of this work.
@@ II Motivation and Examination of Related Work
LLM for Networking: The white paper by Shahid et al. [4] presents the concept of “Large Telecom Model (LTM)” for use cases of telecom networks. Some of the potential use cases they mention include the use of LTM at the network edge, LTM for network optimization, and using LTM for network automation tasks, etc. Similarly, Zhou et al. [5] identify four different areas of telecom networks that may benefit from LLMs. These areas include generation, optimization, classification, and prediction problems in telecom networks. Wu et al. [11] present the NetLLM method to utilize LLMs for three different networking problems, namely, “viewport prediction”, “adaptive bitrate streaming”, and “cluster job scheduling”. The authors extensively evaluate the performance of their proposed framework within these problems.
LLM for Security: Ferrag et al. [6] propose multiple applications of LLMs in cybersecurity, including detection and analysis of threats, incident response, automation of security tasks, etc. The work by Guthula et al. [12] proposes a foundation model for security that takes into account the distinct nature of network traffic. The aim of the authors is to consider the general applicability of the model.
Sędkowski [13] studied the efficacy of LLMs in recognizing potential threats in the network and recommending countermeasures. Their methodology includes using Nmap reports to classify threats using STRIDE threat modeling with three different LLMs. The author concludes that the performance of LLMs in threat detection is comparable to that of humans. The scope of their work is focused on testing the application of AI in threat modeling, as opposed to our aim of studying the suitability of LLMs for network security.
Yang et al. [14] present “ThreatModeling-LLM”, an LLM-based method to perform threat modeling of the banking system using the STRIDE model. Their approach includes various steps to improve the performance of the LLMs. Their scope is limited to banking systems. The main objective of their work is to improve and automate threat modeling using LLMs, instead of investigating the suitability of LLMs for network security. Saha et al. [15] developed “ThreatLens” to perform threat modeling and test plan generation for hardware security verification.
Motivation: After examining the related works, we identify that the suitability of the potential use cases of LLMs needs to be investigated. However, to the best of our knowledge, there are very few existing studies that are actually testing the suitability of LLMs for these use cases. This motivated us to study the suitability of the LLMs in telecom network security using a case study of STRIDE threat modeling.
@@ III Evaluation Methodology for A Case Study of STRIDE Threat Modeling
Our evaluation methodology is shown in Figure 1. We initially select multiple threats and vulnerabilities in the 5G network, along with their baseline STRIDE classifications from the published literature and standards. Then, we use various prompting techniques to perform the LLM-based STRIDE classification of the selected threats. Finally, we compare the STRIDE classification by LLMs to the baseline and evaluate the results. We will explain each step in the following.
@@ III-A Selected 5G Threats and Vulnerabilities
To carry out the evaluation, we select six threats and vulnerabilities on 5G interfaces from our previous research work [10] along with their STRIDE classifications as a baseline; this step is shown with a blue block in Figure 1. As we already mentioned in our previous work [10], we want to clarify that the baseline STRIDE classification of the threats may not be unique. We select these attacks from our previously published study [10] while ensuring that the selected threats cover the six STRIDE categories and span over multiple 5G interfaces. Three of the selected threats are on the N1 interface, because N1 is exposed to the Radio Access Network (RAN) and it faces the largest number of threats [10]. In our earlier work [10], we comprehensively explored and identified the threats and vulnerabilities on the critical 5G network interfaces and categorized them based on the STRIDE threat model. The selected threats, along with their description, are explained below. We use the same threat names in the first row of Table I:
@@ III-B Large-Language Models (LLMs)
We select the following Large Language Models (LLMs) to perform this evaluation: Sonar by Perplexity [20], GPT-4o by OpenAI [21], Claude 3.7 Sonnet by Anthropic [22], Grok-2 by xAI [23], and Gemini 2.5 Pro by Google [24]. We use these LLMs through the pro version of the Perplexity AI platform that we have access to through our University [25]. We are interested in evaluating the suitability of the current LLMs for network security, hence we use the LLMs with base knowledge as it is, without retraining or fine-tuning on any datasets. We perform the STRIDE classification of the six selected threats using these LLMs in order to evaluate the suitability of the LLMs for network security. Figure 1 shows the LLM-based STRIDE classification methodology with green blocks.
@@ III-C LLM Prompts
LLM task, whereas Cyan text refines the output of the LLM
@@
We replace the ‘[NAME_OF_THREAT]’ in these prompts with the name of a specific threat and provide the prompt to the LLM. Then, we record the STRIDE classification of a threat provided by the LLM. We repeat this step for all six selected threats and use the same prompts for each threat as shown in Listings 2 and 3, for each of ZS and FS prompting techniques, respectively (see steps in green blocks in Figure 1).
@@ IV-A LLM-based STRIDE Classification
In this section, we provide the results of evaluating the case study of STRIDE threat modeling of 5G threats. The results of LLM-based STRIDE classification of the six 5G threats are shown in Table I. The first column of the table includes the prompting techniques we employ in our evaluation, while the second column shows the LLMs we select. The rest of the columns in this table present the baseline STRIDE classification along with the results of LLM-based STRIDE classification of the 5G threats. In this table, the white cells with a dot represent a positive baseline value, while an empty white cell represents a negative baseline value. According to this baseline, we categorize the LLM-based STRIDE classifications as True Positive (TP) (dark green cell with a dot ), True Negative (TN) (empty green cell), False Positive (FP) (yellow cell with a dot ), and False Negative (FN) (empty red cell). The coloring scheme represents that the greens (TP and TN) are correct classifications. Yellow (FP) is an incorrect classification, but it is not the worst outcome (over-predicting positive), and red (FN) is an incorrect classification and represents the worst outcome (under-predicting positive). In the following, we present the main insights and observations from these results.
@@ IV-D Performance of LLMs in STRIDE Classification
The results show that LLMs’ performance in our experiments is mostly comparable. Figure 4 illustrates the performance of the LLMs in terms of accuracy, precision, recall, and F1 score using a heatmap chart. This result is averaged across all threats and all prompting techniques for a specific LLM. We observe that GPT-4o, Claude 3.7 Sonnet, and Grok-2 show ‘relatively’ higher accuracy and precision but lower recall compared to the other two LLMs. On the other hand, Sonar and Gemini 2.5 Pro achieved lower precision and higher recall in comparison. Higher recall means that these models capture most of the TP cases and keep FN (the worst outcome) to a minimum. The F1 score indicates that the performance of all the LLMs is comparable for this case study. The maximum accuracy achieved is , which highlights that there are opportunities for improvement across all LLMs, perhaps by fine-tuning for the specific application of STRIDE threat modeling in 5G networks.
Limitations: As we already mentioned in Section III-A, the baseline STRIDE classification of the threats may not be unique. However, we are more interested in investigating the behavior of LLMs and identifying the insights on the LLM-based STRIDE classification, rather than focusing on the accuracy of individual LLM classification.
Challenges: We note that several articles highlight the prominent challenges with the LLMs [28, 29]. The most relevant issues to threat modeling include incorrect predictions and LLM hallucinations, which may result in disregarding required countermeasures or implementing unnecessary security measures. Secondly, the adaptability of LLMs for telecom-specific threats and vulnerabilities. Thirdly, improving LLM inference speed in networks to enable rapid threat modeling of the detected threats.
@@ V Discussion and Conclusion
In this work, we explore and investigate the suitability of LLMs for network threat modeling. To perform this analysis, we select the case study of STRIDE threat modeling to perform LLM-based classification of 5G threats according to the STRIDE threat model. We observe from our evaluation that providing examples to the LLMs using FS prompting improves their performance. We further notice that LLMs may not always consider the threat classifications as a result of the second-order effect. This will limit the threat identification and may eventually result in not identifying all the possible risks associated with a threat. We hope these insights and results of our work are the starting points to encourage research into fine-tuning LLMs on telecom-specific datasets and to enhance their performance in network security tasks. This is particularly important for future ‘AI-native’ networks, where AI needs to detect and identify threats autonomously and with the highest accuracy.